# DiMoDif: Discourse Modality-information Differentiation for Audio-visual Deepfake Detection and Localization

### v1

@@ 3.2 Architecture
Our architecture is illustrated in Figure 2. From each video $\mathfrak{v}=\{v,a\}$, we extract the visual $\mathbf{v}\in\mathbb{R}^{f\times d_{0}}$ and audio $\mathbf{a}\in\mathbb{R}^{f\times d_{0}}$ features corresponding to its visual $v$ and audio $a$ components, denoting with $d_{0}$ the embedding dimension, based on the lip-reading and speech-to-text models proposed in [44]. Under ideal conditions, in real videos, $\mathbf{v}$ and $\mathbf{a}$ contain the same information corresponding to the discourse, while in deepfakes that contain manipulated parts, differentiation is expected. For that reason, the proposed architecture that identifies such patterns is named Discourse Modality-information Differentiation (DiMoDif).
@@
The discourse features, $\mathbf{v}$ and $\mathbf{a}$, are projected to $d$ dimensions using a lightweight 1-D convolutional network with two ReLU-activated layers. Then, we add trainable sequence encoding $S_{v},S_{a}\in\mathbb{R}^{d}$ and employ a separation token $\mathbf{s}\in\mathbb{R}^{d}$ to encode modality information. An additional classification token $\mathbf{c}\in\mathbb{R}^{d}$ is only employed in DFD experiments, and position encoding $P_{j}\in\mathbb{R}^{d}$, with $j=1,\dots,2\cdot f+e$, is applied on all tokens, resulting in tensor $\mathbf{t}=\oplus\{\mathbf{c},\mathbf{v},\mathbf{s},\mathbf{a}\}\in\mathbb{R}^{(2\cdot f+e)\times d}$, with $\oplus$ denoting concatenation, being the input to the Transformer encoder $\mathcal{T}$ [66]. $e$ depends on the task, being equal to 2 in DFD experiments and equal to 1 in TFL experiments where $\mathbf{c}$ is absent. Then, we extract $d$-dimensional feature pyramids corresponding to all $2\cdot f$ visual and audio tokens deriving from all $l$ layers of the Transformer encoder, denoted by $\mathbb{R}^{(2\cdot f)\times l\times d}\ni\mathbf{z}\subset\mathcal{T}(\mathbf{t})$. To do so, $\mathcal{T}$ processes $\mathbf{t}$ as:
$\tilde{\mathbf{z}}_{j}$ $=\text{MSA}(\text{LN}(\mathbf{z}_{j-1}))$ (2)
$\mathbf{z}_{j}$ $=\text{MLP}(\text{LN}(\tilde{\mathbf{z}}_{j}))+\tilde{\mathbf{z}}_{j}$ (3)
@@
where $j=1,\dots,l$ denotes the layer index, $\mathbf{z}_{0}=\mathbf{t}$ is the input, MSA denotes Multi-head Softmax Attention [66] with $r$ number of heads, LN denotes Layer Normalization [37], and MLP denotes a Multi-Layer Perceptron with internal dimensionality $d\cdot u$. Then, $\mathbf{z}=\oplus\{\mathbf{z}_{j}\}_{j=1}^{l}$, omitting the tokens that correspond to $\mathbf{c}$ and $\mathbf{s}$. Given that short-term inter-modality inconsistency detection is key to our tasks, local attention is preferable in contrast to global; besides, it is computationally more efficient. Thus, we limit the Transformer’s attention to a local window $[j-q/2,j+q/2]$ of $q$ tokens around each token $j$.
Figure 2: The DiMoDif architecture.
@@
Finally, in DFD experiments, predictions $\hat{y}$ are made by a feed-forward classification head $\mathcal{Q}$ that takes as input the classification tokens of $\mathbf{z}$, denoted by $\mathbf{z}^{\prime}\in\mathbb{R}^{l\times d}$, and consists of three linear layers, the first two of which are followed by Layer Normalization (LN) and ReLU activation:
$\mathbf{z}^{\prime\prime}$ $=\text{ReLU}\big{(}\text{LN}(\mathbf{z}^{\prime}\cdot\mathrm{W}_{1}+\mathrm{b}_{1})\big{)}$ (4)
$\mathbf{z}^{\prime\prime\prime}$ $=\text{ReLU}\big{(}\text{LN}(\mathbf{z}^{\prime\prime}\cdot\mathrm{W}_{2}+\mathrm{b}_{2})\big{)}$ (5)
$\hat{y}$ $=\mathbf{z}^{\prime\prime\prime}\cdot\mathrm{W}_{3}+\mathrm{b}_{3}$ (6)
@@ 3.3 Objective Function
For the DFD task, we consider the binary cross-entropy loss [24], directly optimizing the real vs. fake objective. For the TFL task, we propose a composite loss as the combination of three loss functions, namely (1) the focal loss [43] to optimize the classification objective, while accounting for class imbalance being prevalent in each video’s frames, (2) the DIoU loss [79] to optimize the regression objective by maximizing the overlap between predicted and ground truth fake time intervals, and (3) the smooth L1 loss [23] to minimize the distance between predicted and actual boundaries of fake time intervals. Detailed mathematical expressions explaining the computation of both loss functions can be found in the Appendix A.
@@ 4.1 Datasets
We consider FakeAVCeleb [34], LAV-DF [10], and AV-Deepfake1M [8] datasets for training and evaluation. Appendix B provides details on class and split sizes. FakeAVCeleb contains 21K fake and 500 real samples, thus for balance we added 20K more from VoxCeleb2 dataset [15] including 10K of male and 10K of female speakers of several ethnicities talking in English [^4] Language identification conducted with https://github.com/speechbrain/speechbrain., reaching a total of 41,544 samples randomly split based on a 80-10-10 identity separation. LAV-DF has been released with standard splits of 78K training, 31K validation, and 26K test samples. AV-Deepfake1M has 746K training, 57K validation, and 343K test samples the metadata of which have not been released. We also evaluate DiMoDif on the test set of DFDC [20] and a randomly selected 6K-sized sample from KoDF [36] proportionally spanning all synthesis methods and real samples, to assess generalization.
@@ 4.2 Evaluation
For DFD, we use accuracy (ACC), Average Precision (AP), and area under ROC curve (AUC). For TFL, we use Average Precision at $p$ (AP@$p$), and Average Recall at $n$ (AR@$n$). We report metrics per dataset in accordance to previous works for comparability. On AV-Deepfake1M we evaluate through Codabench [^5] https://deepfakes1m.github.io/evaluation. Competing methods are audio-visual deepfake detectors, face manipulation detectors, and temporal action localization models, discussed in Section 2.
@@ 4.3 Implementation Details
We train DiMoDif for 100 epochs with early stopping patience 10 and checkpoint based on validation metrics, AUC for DFD and the sum of AP@$p$ and AR@$n$ for TFL. Batch size is set to 64, initial learning rate to 0.001 with ReduceLROnPlateau scheduler [^6] https://pytorch.org/docs/stable/generated/torch.optim.lr_scheduler.ReduceLROnPlateau.html and Adam optimizer, local attention window size $q$ to 15, and focal loss’ hyperparameters $\alpha=0.98$ and $\gamma=2$. Also, we set a maximum sequence length $f$ to 600 and zero-pad the smaller. We consider a grid for $\mathcal{T}$’s size, with $(d,r,u)\in\{(32,2,1),(64,2,1),(64,4,1),(128,4,2),(256,8,2)\}$ and number of layers $l\in\{1,3,5\}$. Grid search for AV-Deepfake1M is performed with 200K training samples to reduce resource requirements, but retrain the best configuration on the whole set. We spent $\sim$3900 GPU hours for training on NVIDIA GeForce RTX 3090 Ti GPUs.
@@ 5 Results
5.1 Temporal Forgery Localization
Method Modality AP@0.5 AP@0.75 AP@0.95 AR@100 AR@50 AR@20 AR@10
MDS [14] $\mathcal{AV}$ 12.8 1.6 0.0 37.9 36.7 34.4 32.2
AGT [50] $\mathcal{V}$ 17.9 9.4 0.1 43.2 34.2 24.6 16.7
BMN [42] $\mathcal{V}$ 24.0 7.6 0.1 53.3 41.2 31.6 26.9
BMN (I3D) [42] $\mathcal{V}$ 10.6 1.7 0.0 48.5 44.4 37.1 31.6
AVFusion [5] $\mathcal{AV}$ 65.4 23.9 0.1 63.0 59.3 54.8 52.1
ActionFormer [74] $\mathcal{V}$ 95.3 90.2 23.7 88.4 89.6 90.3 90.4
BA-TFD [10] $\mathcal{AV}$ 76.9 38.5 0.3 66.9 64.1 60.8 58.4
BA-TFD+ [9] $\mathcal{AV}$ 96.3 85.0 4.4 81.6 80.5 79.4 78.8
UMMAFormer [76] $\mathcal{AV}$ 98.8 95.5 37.6 92.4 92.5 92.5 92.1
DiMoDif (ours) $\mathcal{AV}$ 95.5 87.9 20.6 94.2 93.7 92.7 91.4
TriDet+InternVideo [61, 68] $\mathcal{V}$ 29.66 09.02 00.79 00.09 24.08 24.06 23.96 23.50 22.55
ActionFormer+InternVideo [74, 68] $\mathcal{V}$ 36.08 12.01 01.23 00.16 27.11 27.08 27.00 26.60 25.80
Table 3: Temporal forgery localization results on AV-Deepfake1M [8]. Modality denotes the model’s input type with $\mathcal{V}$ being visual and $\mathcal{A}$ audio. Bold indicates best and underline second to best performance.
Tables 2 and 3 present the performance of our method in comparison to state-of-the-art, in terms of Temporal Forgery Localization (TFL), on LAV-DF [10] and AV-Deepfake1M [8] datasets, respectively. On LAV-DF, DiMoDif outperforms all competitive approaches wrt AR@{100,50,20}, is second to best wrt AR@10, and third wrt AP@{0.5,0.75,0.95}. UMMAFormer [76], BA-TFD+ [9] and ActionFormer [74] exhibit similar performance, while the rest of the models have lower localization abilities. On AV-Deepfake1M, DiMoDif outperforms all competitive approaches wrt all metrics exhibiting significant performance increase. The proposed method’s AR is almost double compared with the second to best, while it exhibits a notable absolute increase of +35.29% and +47.88% in AP at 0.5 and 0.75, respectively.
@@ 5.1 Temporal Forgery Localization
DiMoDif (ours) $\mathcal{AV}$ 99.6
@@
DiMoDif (ours) $\mathcal{AV}$ 96.3 83.3*
Table 5: Deepfake detection results on AV-Deepfake1M [8]. Modality denotes the model’s input type with $\mathcal{V}$ being visual and $\mathcal{A}$ audio. Bold indicates best and underline second to best performance. *Validation set accuracy is reported as Codabench only provides AUC score.
@@
Method Modality Pre-training RVFA FVRA-WL FVFA-WL FVFA-FS FVFA-GAN AVG-FV
@@
DiMoDif (ours) $\mathcal{AV}$ 84.5 97.1 99.0 99.9 99.7 99.9 99.5 99.9 99.6 99.9 99.4 99.9
Table 6: Deepfake detection results on FakeAVCeleb [34]. Modality denotes the model’s input type with $\mathcal{V}$ being visual and $\mathcal{A}$ audio. Bold indicates best and underline second to best performance. Supervised methods that use pre-training are fine-tuned on FakeAVCeleb, while unsupervised methods are not trained with labels and fake examples.

### v2

@@ 3.2 Discourse-related Feature Extraction (DiFE)
Discourse modality-information differentiation (DiMoDif) is learned by utilizing speech features extracted with disjoint-only encoding of visual and audio streams. Otherwise, e.g., by utilizing Audio-Visual Speech Recognition (AVSR) encoders, information missing from one modality due to deepfake artifacts is retrieved from the other resulting in indistinguishable cross-modal differences between real and fake samples. From each video $\mathfrak{v}=\{v,a\}$, we extract the visual $\mathbf{v}\in\mathbb{R}^{f\times d_{0}}$ and audio $\mathbf{a}\in\mathbb{R}^{f\times d_{0}}$ features corresponding to its visual $v$ and audio $a$ components, denoting with $d_{0}$ the embedding dimension, based on the distinct VSR and ASR models proposed in [46]. Then, $\mathbf{v}$ and $\mathbf{a}$ are projected to $d$ dimensions using a lightweight 1-D convolutional network with two ReLU-activated layers. Under ideal conditions, in real videos, $\mathbf{v}$ and $\mathbf{a}$ contain the same information corresponding to the discourse, while in deepfakes that contain manipulated parts, differentiation is expected due to artifacts.
3.3 Modality-information Differentiation (MiD)
@@
Trainable sequence encoding $S_{v},S_{a}\in\mathbb{R}^{d}$ is added to $\mathbf{v}$ and $\mathbf{a}$, while a separation token $\mathbf{s}\in\mathbb{R}^{d}$ is employed to encode modality information. An additional classification token $\mathbf{c}\in\mathbb{R}^{d}$ is only employed in DFD experiments, and position encoding $P_{\tau}\in\mathbb{R}^{d}$, with $\tau=1,\dots,2\cdot f+e$, is applied on all tokens, resulting in tensor $\mathbf{t}=\oplus\{\mathbf{c},\mathbf{v},\mathbf{s},\mathbf{a}\}\in\mathbb{R}^{(2\cdot f+e)\times d}$, with $\oplus$ denoting concatenation, being the input to the Transformer encoder $\mathcal{T}$ [70]. $e$ depends on the task, being equal to 2 in DFD experiments and equal to 1 in TFL experiments where $\mathbf{c}$ is absent. Given that short-term inter-modality inconsistency detection is key to our tasks, multimodal local attention is preferable in contrast to global; besides, it is computationally more efficient. Thus, we limit the Transformer’s attention to a multimodal local window of $2\cdot q$ tokens $\tau\in[1+\phi-\left\lfloor q/2\right\rfloor,1+\phi+\left\lfloor q/2\right\rfloor]\cup[f+2+\phi-\left\lfloor q/2\right\rfloor,f+2+\phi+\left\lfloor q/2\right\rfloor]$ for frames $\phi$, while $\mathbf{c}$ and $\mathbf{s}$ attend to all tokens. Then, we extract $d$-dimensional feature pyramids corresponding to $\mathbf{c}$ (if applicable), $\mathbf{v}$, and $\mathbf{a}$ tokens deriving from the $l$ layers of the Transformer $\mathcal{T}$, denoted by $\Phi\subset\mathcal{T}(\mathbf{t},q)$, with $\Phi\in\mathbb{R}^{(2\cdot f+e-1)\times l\times d}$. To do so, $\mathcal{T}$ processes $\mathbf{t}$ as:
$\tilde{\mathbf{z}}_{\lambda}$ $=\text{MSA}(\text{LN}(\mathbf{z}_{\lambda-1}))$ (2)
$\mathbf{z}_{\lambda}$ $=\text{MLP}(\text{LN}(\tilde{\mathbf{z}}_{\lambda}))+\tilde{\mathbf{z}}_{\lambda}$ (3)
@@
where $\lambda=1,\dots,l$ denotes the layer index, $\mathbf{z}_{0}=\mathbf{t}$ is the input, MSA denotes Multi-head Softmax Attention [70] with $r$ number of heads, LN denotes Layer Normalization [39], and MLP denotes a Multi-Layer Perceptron with internal dimensionality $d\cdot u$. Then, $\Phi=\oplus\{\mathbf{z}_{\lambda}\}_{\lambda=1}^{l}$, omitting the separation token $\mathbf{s}$.
3.4 Classification and Regression Heads
@@
In DFD experiments, predictions $\hat{y}$ are made by a feed-forward classification head $\mathcal{Q}$ that takes as input the classification tokens of $\Phi$, denoted by $\Psi\in\mathbb{R}^{l\times d}$, and consists of three linear layers, the first two of which are followed by Layer Normalization (LN) and ReLU activation:
$\dot{\Psi}$ $=\text{ReLU}\big{(}\text{LN}(\Psi\cdot\mathrm{W}_{1}+\mathrm{b}_{1})\big{)}$ (4)
$\ddot{\Psi}$ $=\text{ReLU}\big{(}\text{LN}(\dot{\Psi}\cdot\mathrm{W}_{2}+\mathrm{b}_{2})\big{)}$ (5)
$\hat{y}$ $=\ddot{\Psi}\cdot\mathrm{W}_{3}+\mathrm{b}_{3}$ (6)
@@ 3.5 Objective Function
For the DFD task, we consider the binary cross-entropy loss $\mathfrak{L}_{ce}$ [25], optimizing the real vs. fake objective. For the TFL task, we propose a composite loss as the combination of three loss functions, namely (1) the focal loss $\mathfrak{L}_{f}$ [45] to optimize the classification objective, while accounting for class imbalance being prevalent in each video’s frames, (2) the DIoU loss $\mathfrak{L}_{d}$ [83] to optimize the regression objective by maximizing the overlap between predicted and ground truth fake time intervals, and (3) the smooth L1 loss $\mathfrak{L}_{1}^{s}$ [24] to minimize the distance between predicted and actual boundaries of fake time intervals. Detailed mathematical expressions of loss function computation can be found in the supplementary material.
@@ 4.1 Datasets
We consider FakeAVCeleb [36], LAV-DF [10], and AV-Deepfake1M [8] datasets for training and evaluation. FakeAVCeleb contains 21K fake and 500 real samples which we split in 70% for training and 30% for testing, following [75, 54] [^5] To balance classes during training only, we add real samples from VoxCeleb2 [16] including equal number of male and female speakers of several ethnicities talking in English. Language identification via https://github.com/speechbrain/speechbrain., while we keep 5% of the training set for validation. LAV-DF has been released with standard splits of 78K training, 31K validation, and 26K test samples. AV-Deepfake1M has 746K training, 57K validation, and 343K test samples; the metadata of the latter have not been released. Supplementary material provides further details on class and split sizes.
@@ 4.2 Evaluation
For DFD, we use accuracy (ACC), Average Precision (AP), and area under ROC curve (AUC). For TFL, we use Average Precision at $p$ (AP@$p$), and Average Recall at $n$ (AR@$n$). We report metrics per dataset in accordance to previous works for comparability. On AV-Deepfake1M we evaluate through Codabench platform [^6] https://deepfakes1m.github.io/evaluation. Competing methods are audio-visual deepfake & face manipulation detectors, and temporal action localization models (cf. Section 2). Also for DFD the naive thresholding on $\mathtt{d_{L}}$ approach (cf. fig. 2).
@@ 4.3 Implementation Details
We train DiMoDif for 100 epochs with early stopping patience 10 and checkpoint based on validation metrics, AUC for DFD and the sum of AP@$p$ and AR@$n$ for TFL. Batch size is set to 64, initial learning rate to 0.001 (reduced on plateau) with Adam optimizer, local attention window size $q$ to 15, and focal loss’ hyperparameters $\alpha=0.98$ and $\gamma=2$. We set a maximum sequence length $f$ to 600 and zero-pad the smaller. We consider a grid for $\mathcal{T}$’s size, with $(d,r,u)\in\{(32,2,1),(64,2,1),(64,4,1),(128,4,2),(256,8,2)\}$ and number of layers $l\in\{1,3,5\}$. Grid search for AV-Deepfake1M is performed with 200K training samples to reduce resource requirements, but retrain the best configuration on the whole set. We spent $\sim$3900 GPU hours for training on NVIDIA GeForce RTX 3090 Ti GPUs.
@@ 5 Results
5.1 Ablation Study
Other speech-related pre-trained models are equivalent for use as backbones in DiMoDif, e.g., AV-Hubert [63] (cf. section 3.2). To study their impact, we train DiMoDif on FakeAVCeleb and LAV-DF w/ AV-Hubert features and present the results in Table 2. This variant performs slightly worse on FakeAVCeleb and LAV-DF (DFD), better wrt AP@0.5 and worse wrt AR@100 on LAV-DF (TFL). Henceforth, we opt for the better performing and more intuitive speech prediction variant (Ma 2022) [46]. Figure 4 presents ablations on AV-Deepfake1M [^7] We report validation performance as conducting the ablations through Codabench would be impractical.. Figure 4(a) indicates that a small window size $q=15$ for Transformer’s $\mathcal{T}$ local attention is optimal, especially in contrast to attending on all tokens determined by $q=0$. Also, the feature pyramid scheme provides a small performance increase as indicated by Figure 4(b). The size of $\mathcal{T}$ is important for achieving maximum performance as shown in Figure 4(c), both in terms of number of layers $l$ and layer size determined by input dimension $d$, number of attention heads $r$, and internal dimension $d\cdot u$: the larger the model the higher its performance. Supplementary material provides further TFL task ablations on AV-Deepfake1M wrt hyperparameter $\alpha$ and learning rate scheduler, the same analysis on LAV-DF, and a pertinent DFD task ablation analysis.
FAVC L-DF (DFD) L-DF (TFL)
AV-Hubert [63] 99.98, 99.35* 99.79, 99.78* 98.00+, 93.00
(Ma 2022) [46] 99.99, 99.71* 99.94, 99.84* 95.49+, 94.17
Table 2: Speech-related pre-trained backbone ablation analysis. AP, *AUC, +AP@0.5, AR@100.
(a) AV-Deepfake1M: Window size $q$
(b) AV-Deepfake1M: Feature pyramid (FP) use
(c) AV-Deepfake1M: $\mathcal{T}$ size; $l$ $(d,r,d\cdot u)$
Figure 4: Ablation and hyperparameter tuning analysis.
5.2 Deepfake Detection
Tables 5, 5 and 5, present comparative results in terms of video-level Deepfake Detection (DFD) on FakeAVCeleb [36], LAV-DF [10], and AV-Deepfake1M [8]. On FakeAVCeleb, DiMoDif outperforms competitive methods achieving +0.8 ACC, and +0.6 AUC. On LAV-DF, DiMoDif performs slightly better than UMMAFormer [80], achieving 99.84 AUC. On AV-Deepfake1M dataset, DiMoDif outperforms the state-of-the-art by a significant absolute performance increase of +30.5 AUC [^8] Codabench does not provide ACC for the test samples, thus we report the validation accuracy, which, even though is not comparable, still provides a sense on the level of uncalibrated threshold performance increase.. Table 6 presents cross-manipulation performance on FakeAVCeleb in which DiMoDif achieves 100.0 AVG-FV AP yet 66.4 AP on RVFA. Notably, DiMoDif presents perfect detection rate in face-swaps (100.0 AUC). Mouth in face-swaps moves occasionally unnaturally yielding inaccurate lip-reading predictions, while audio is unaffected yielding correct speech-to-text predictions. DiMoDif translates this inconsistency into a deepfake detection. The naive classifier achieves high performance on FakeAVCeleb but moderate performance on LAV-DF and AV-Deepfake1M; however, it outperforms several methods. Finally, challenging real samples with $\mathtt{d_{L}}$ over 0.8 (cf. section 1), are scored by DiMoDif with an average of 0.19, and 0.16 on LAV-DF, and AV-Deepfake1M, respectively, attesting to DiMoDif’s advantage wrt naive distance-thresholding.
Method Modality ACC AUC
naive $\mathcal{AV}$ 97.7 93.3
Xception [60] $\mathcal{V}$ 67.9 70.5
LipForensics [30] $\mathcal{V}$ 80.1 82.4
FTCN [82] $\mathcal{V}$ 64.9 84.0
MDS [15] $\mathcal{AV}$ 82.8 86.5
AVoiD-DF [75] $\mathcal{AV}$ 83.7 89.2
ART-AVDF [73] $\mathcal{AV}$ 96.4 98.2
AVFF [54] $\mathcal{AV}$ 98.6 99.1
DiMoDif (ours) $\mathcal{AV}$ 99.4 99.7
Table 3: In-dataset perfromance on FakeAVCeleb [36]. Modality denotes the model’s input type with $\mathcal{V}$ being visual and $\mathcal{A}$ audio. Bold indicates best and underline second to best performance.
@@ 5.2 Deepfake Detection
DiMoDif (ours) $\mathcal{AV}$ 99.84
@@
DiMoDif (ours) $\mathcal{AV}$ 96.3 96.3*
Table 5: Deepfake detection results on AV-Deepfake1M [8]. Modality denotes the model’s input type with $\mathcal{V}$ being visual and $\mathcal{A}$ audio. Bold indicates best and underline second to best performance. *Computed on validation set.
@@
Method Modality Pre-training RVFA FVRA-WL FVFA-FS FVFA-GAN FVFA-WL AVG-FV
@@
AVFF [54] $\mathcal{AV}$ LRS3 [3] 93.3 92.4 94.8 98.2 100. 100. 99.9 100. 99.4 99.8 98.5 99.5
DiMoDif (ours) $\mathcal{AV}$ 66.4 51.6 100. 99.8 100. 100. 100. 100. 100. 100. 100. 99.9
Table 6: Cross-manipulation performance on FakeAVCeleb [36]. Modality denotes the model’s input type with $\mathcal{V}$ being visual and $\mathcal{A}$ audio. Bold indicates best and underline second to best performance. Supervised methods that use pre-training are fine-tuned on FakeAVCeleb, while unsupervised methods are not trained with labels and fake examples.
5.3 Temporal Forgery Localization
Method Modality AP@0.5 AP@0.75 AP@0.95 AR@100 AR@50 AR@20 AR@10
MDS [15] $\mathcal{AV}$ 12.8 1.6 0.0 37.9 36.7 34.4 32.2
AGT [51] $\mathcal{V}$ 17.9 9.4 0.1 43.2 34.2 24.6 16.7
BMN (I3D) [44] $\mathcal{V}$ 10.6 1.7 0.0 48.5 44.4 37.1 31.6
AVFusion [5] $\mathcal{AV}$ 65.4 23.9 0.1 63.0 59.3 54.8 52.1
ActionFormer [78] $\mathcal{V}$ 95.3 90.2 23.7 88.4 89.6 90.3 90.4
BA-TFD [10] $\mathcal{AV}$ 76.9 38.5 0.3 66.9 64.1 60.8 58.4
BA-TFD+ [9] $\mathcal{AV}$ 96.3 85.0 4.4 81.6 80.5 79.4 78.8
UMMAFormer [80] $\mathcal{AV}$ 98.8 95.5 37.6 92.4 92.5 92.5 92.1
MMMS-BA [34] $\mathcal{AV}$ 97.5 95.2 39.0 94.0 93.4 95.9 89.4
DiMoDif (ours) $\mathcal{AV}$ 95.5 87.9 20.6 94.2 93.7 92.7 91.4
PyAnnote (Zero-Shot) [56] $\mathcal{A}$ 00.03 00.00 00.00 00.00 00.67 00.67 00.67 00.67 00.67
MMMS-BA [34] $\mathcal{AV}$ 62.75* 35.87* - 18.37* 57.49* - 55.94* 54.28* -
Table 8: Temporal forgery localization results on AV-Deepfake1M [8]. Modality denotes the model’s input type with $\mathcal{V}$ being visual and $\mathcal{A}$ audio. Bold indicates best and underline second to best performance. *Computed on validation set.
Tables 7 and 8 present the performance of DiMoDif in comparison to state-of-the-art, in terms of Temporal Forgery Localization (TFL), on LAV-DF [10] and AV-Deepfake1M [8], respectively. On LAV-DF, DiMoDif outperforms all competitive approaches wrt AR@{100,50}, is second to best wrt AR@{20,10}, and performs below state-of-the-art wrt AP@{0.5,0.75,0.95}. MMMS-BA [34], UMMAFormer [80], BA-TFD+ [9] and ActionFormer [78] exhibit similar performance, while the remaining models have lower localization ability. On AV-Deepfake1M, DiMoDif outperforms all competitive approaches wrt all metrics exhibiting significant performance increase. The proposed method’s AR is almost double compared with the second to best, while it exhibits a notable absolute increase of +35.29% and +47.88% in AP at 0.5 and 0.75, respectively.
5.4 Interpretability
Figure 5: Fake video sample (manipulated in red). Avg. and std. (across layers $\lambda$) of cross-modal similarity and frame-level fake probability shown in blue and purple, respectively.
DiMoDif is inherently interpretable as its frame-level cross-modal representations’ similarity reflects the level of congruity between visual and audio speech. Figure 5 illustrates a fake video example from LAV-DF [^9] LAV-DF/test/000010.mp4 with one small manipulated part during which cross-modal similarity is significantly reduced indicating audio-visual incongruity while during real parts similarity is close to 1.0. Fake frames are perfectly identified by DiMoDif with $\max_{m}(\hat{a}^{m}_{\phi})>>0.5$. More real and fake examples are provided in suppl. material.
5.5 Robustness
Figure 6: Robustness to visual (Gaussian blur) and audio (Gaussian noise) distortions.
Figure 6 illustrates DiMoDif’s robustness to visual (Gaussian blurring) and audio (Gaussian noise) distortions at 5 levels of intensity. The analysis is conducted on FakeAVCeleb test set, following the evaluation protocol of [54, 23, 30] [^10] Perturbations’ implementation directly sourced from https://github.com/EndlessSora/DeeperForensics-1.0.. DiMoDif is not affected by the audio perturbations, outperforming AVFF [^11] Code unavailable. Value approximations from [54]’s supplementary material. [54] by approx. + 6.5 AUC at intensity level 5. Robustness to visual perturbations is gradually reduced to 96.5 AUC, approx. +25.0 AUC compared to AVFF. Supplementary material shows that DiMoDif is also more robust than AVFF under saturation, contrast, block-wise, and JPEG compression visual perturbations, yet less robust under Gaussian noise and video compression. Also, DiMoDif is more robust than AVFF under pitch shift, reverberence, and audio compression audio perturbations, achieving an unprecedented 99.3 AUC on average at intensity level 5 indicating the high robustness of speech prediction features on such distortions.
5.6 Generalization
This paper is the first to conduct generalization experiments on the considered datasets. Table 12 presents DiMoDif’s cross-dataset performance on DFD. The reported performance, although high, indicates that the training data affect generalization due to scale, fake part duration, and generator quality. A small pool of low quality easier fake data (FakeAVCeleb) makes the model prone to overfitting reducing its generalization ability to partially manipulated data generated through more complex approaches (LAV-DF, AV-Deepfake1M). In contrast, large pools of high-quality and harder deepfakes (LAV-DF, AV-Deepfake1M) result in richer and more robust representations, and less overfitting. Supplementary material provides further generalization evaluations.